# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

### v1

@@ 7.2.5 Transfer relevance
To provide an overview of how much event data of the target context should be available to train a model outperforming our transferred models (not fine-tuned), Figure 8 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the glove-wiki-gigaword-100 (see Section 7.2.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.2.4) for time encoding.
@@ 8.1 Theoretical implications
Second, our results in the inter- and intra-organizational use cases indicate that pre-trained embedding models can be a better choice to encode the activities for TL in PPM than using a sparse encoding technique, such as one-hot encoding. In addition, depending on the domain (event data) in the source and target contexts, different pre-trained embedding models lead to the best prediction performance of the transferred model in the target context. One possible explanation for this is that there is a correlation between the data used to pre-train the embedding models and the activity attribute values in the event data of the source and target contexts.
Third, we show how embedding models can facilitate TL in PPM. By using these models, process activities from different business processes of the same type, having different names but similar semantic meanings, can be encoded accordingly. This is possible with embedding models because they are trained in a self-supervised manner with auxiliary tasks (e.g., predicting a word in a sentence given the remaining words) to learn the representation of words while capturing their semantics.
Moreover, in our experiments, we observed that two important factors for the transfer of knowledge from one process to another are i) the prediction target and ii) the process information considered for the transfer. Regarding the prediction target, a process-agnostic binary outcome prediction target, like the in-time prediction, is easier for the transfer than a process-specific outcome prediction or the next activity prediction. This is because the latter prediction targets are conceptually more different in the source and the target context. Concerning the considered process information, a transfer that relies solely on the control flow is typically easier than a transfer that relies on the control flow and additional time or data attributes. This is because the consideration of each additional attribute in the event data of the source and target contexts, with its values and distribution, increases the complexity. Nevertheless, to show that information beyond the control flow can be transferred from source to target contexts, we additionally considered the timestamp attribute in the intra- and inter-organization use cases of our evaluation and extracted the time feature duration since start as part of the relative cross-domain encoding.
@@ 8.2 Practical implications
Our work also has implications for practice. With a TL-based PPM technique like ours, the adoption of PPM may be fostered in practice. This is specifically beneficial for organizations or areas in organizations without suitable resources, such as computational power, skills or training of employees, or event data for model training. These organizations can use the base model and further resources from another organization that already uses PPM successfully for similar business processes. This allows them to predict, for example, process outcomes or throughput times and, therefore, mitigate risks and decrease costs.
Furthermore, the proposed technique is scalable to incorporate the event data of multiple processes of the same type in the source and target contexts from a conceptual perspective. In this research, we used one event log as a source domain to train our base model, and one event log as a target domain to apply the base model. However, it is conceivable that multiple event logs can be used as a domain in the source context for training our base model. Similarly, multiple event logs can be used as a domain in the target context to apply our base model. To scale our technique in such a way, concepts from the field of federated learning can be incorporated (Verma et al. 2019). As our technique uses an embedding model for activity encoding, scaling is easier than, for example, with a fixed encoding strategy like one-hot encoding, where each activity included in the event data of the source and target contexts has a specific position in the vector.
@@ 8.3 Future research
There are four promising directions for future research. First, fine-tuning of the embedding models used in our technique for feature-based TL can be investigated. In doing so, we suspect an improvement in the prediction performance of the model used in the target context. As the source context can comprise multiple business processes and their event data, certain strategies for fine-tuning can be realized to adapt the embedding models to multiple similar processes in the same context in a controlled way. For example, such a strategy could be to randomly take 1,000 traces of each business process’s event log to ensure a fair fine-tuning across multiple business processes with different amounts of event data.
@@
Second, the transfer of information about additional event log dimensions can be investigated. In this paper, we have considered information on the control flow and the time perspectives in the design of our technique. However, information on additional attributes available in the event data, such as data attributes, might be valuable for the transfer in PPM. For this purpose, approaches from the research stream of multidimensional process representation creation and learning in BPM (Weinzierl et al. 2024) can be helpful to find a representation of multiple event log attributes, facilitating the process knowledge transfer from source to target contexts. Closely connected to the information used and its representation lies the question of which information is necessary for the knowledge transfer. This question can be investigated with explainable artificial intelligence (XAI) approaches, which are common in PPM (Stierle et al. 2021).
@@
Third, the application of TL in the context of real business processes of different types can be investigated. In the case of those business processes, the complexity of TL is further increased as the business processes are more different from a structural and semantic perspective. For example, most of the activities, as well as additional data attributes in the event data of a business process in the source context, can differ from the event data of a business process in the target context.
@@
Fourth, the building of process-specific foundation models that are universally usable for various data-driven BPM tasks along the BPM lifecycle (Rizk et al. 2023; Buss et al. 2024) is another promising avenue for future research. For this, existing foundation models for (multivariate) time series can serve as a starting point for model development (e.g., Ye et al. 2024), while available benchmark event logs and other process-relevant resources (e.g., fundamental BPM books, scientific BPM papers, or additional descriptions of event logs) can be used as a data basis for initial model training. As foundational models can support multiple modalities, types of process data going beyond text-transformed event data (e.g., process model collections or process-attached images, audio tracks, or video sequences) are additionally conceivable for model fine-tuning. To fine-tune a foundation model with event data from various organizations, the concept of federated learning can be essential to overcome data privacy restrictions (Verma et al. 2019).
@@ 9 Conclusion
In this paper, we proposed a TL-based technique for PPM, allowing organizations without suitable event data or other relevant resources to generate predictions in running business processes for effective decision support. For this purpose, we applied computational DSR (Rai et al. 2017) and designed an artifact in the form of a TL-based technique for PPM, consisting of three phases: i) initial model building on source, ii) transfer base model from source to target, and iii) online application of model to target. By instantiating our technique in the context of intra- and inter-organizational use cases, including event logs for ITSM processes from two organizations, we demonstrated and evaluated its prediction performance. Our results suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context.
@@ Appendix A Prediction performance for complete model compared to model sub-variants
Table 12 shows that the complete model of our proposed TL-based technique clearly outperforms its sub-variants in terms of prediction performance when it is trained on $L_{S}$ and tested on $L_{T}$ for the inter-organizational use case. This confirms the importance of using both components in the prediction model for an effective process knowledge transfer from the source to the target contexts.
@@ Appendix B Prediction performance for different prefix encoding approaches
We applied the index-based encoding, the last-3-state encoding, and the aggregation encoding to our proposed TL-based PPM technique for the inter-organizational use case. We also applied these three encoding approaches to two traditional PPM techniques, using LSTM and XGBoost models. We selected these two traditional PPM techniques because the first uses the same LSTM model as in our proposed technique but not in an TL-setting, and the second uses an XGBoost model, which performed relatively well in the first evaluation of our inter-case organizational use case (see Section 7.2.1.
@@
Table 14: Prediction performance for our TL-based PPM technique that encodes activity and timestamp information about a trace separately compared to a baseline technique that encodes activity and timestamp information about a trace jointly. The baseline technique was tested using the four pre-trained contextualized embeddings of our setup for the inter-organizational use case (average and standard deviation over five runs).

### v2

@@ 7.2.5 Transfer relevance
To provide an overview of how much event data of the target context should be available to train a model outperforming our transferred models (not fine-tuned), Figure 8 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of traces using the glove-wiki-gigaword-100 embedding model (see Section 7.2.2) for activity encoding and the relative cross-domain approach (see Section 7.2.4) for time encoding.
@@ 8.1 Theoretical implications
Second, our results in the inter- and intra-organizational use cases indicate that pre-trained embedding models can be a better choice to encode the activities for TL in PPM than using a sparse encoding technique, such as one-hot encoding. In addition, depending on the domain (event data) in the source and target context, different pre-trained embedding models lead to the best prediction performance of the transferred model in the target context. One possible explanation for this is that there is a correlation between the data used to pre-train the embedding models and the activity attribute values in the event data of the source and target contexts.
Third, we show how embedding models can facilitate TL in PPM. These models enable the encoding of process activities from different business processes of the same type that have different names but similar semantic meanings. This is possible with embedding models because they are trained in a self-supervised manner with auxiliary tasks (e.g., predicting a word in a sentence given the remaining words) to learn the representation of words while capturing their semantics.
Moreover, in our experiments, we observed that two important factors for the transfer of knowledge from one process to another are i) the prediction target and ii) the process information considered for the transfer. Regarding the prediction target, a process-agnostic binary outcome prediction target, like the in-time prediction, is easier for the transfer than a process-specific outcome prediction or the next activity prediction. This is because the latter prediction targets differ more conceptually in the source and target contexts. Concerning the considered process information, a transfer that relies solely on the control flow is typically simpler than a transfer that relies on the control flow and additional time or data attributes. This is because the consideration of each additional attribute in the event data of the source and target contexts, with its values and distribution, increases the complexity. Nevertheless, to show that information beyond the control flow can be transferred from source to target contexts, we additionally considered the timestamp attribute in the intra- and inter-organization use cases of our evaluation and extracted the time feature duration since start as part of the relative cross-domain encoding.
@@ 8.2 Practical implications
Our work also has implications for practice. With a TL-based PPM technique like ours, the adoption of PPM can be encouraged in practice. This is specifically beneficial for organizations or areas in organizations without suitable resources, such as computational power, skills or training of employees, or event data for model training. These organizations can use the base model and further resources from another organization that already uses PPM successfully for similar business processes. This enables them to predict, for example, process outcomes or throughput times, and to use these predictions to mitigate risks and decrease costs.
Furthermore, the proposed technique is scalable to incorporate the event data of multiple processes of the same type in the source and target contexts from a conceptual perspective. In this research, we used one event log as a source domain to train our base model, and one event log as a target domain to apply the base model. However, it is conceivable that multiple event logs can be used as a domain in the source context for training our base model. Similarly, multiple event logs can be used as different domains in the target context to which our base model is applied. To scale our technique in such a way, concepts from the field of federated learning can be incorporated (Verma et al. 2019). As our technique uses an embedding model for activity encoding, scaling is easier than, for example, with a fixed encoding strategy like one-hot encoding, where each activity included in the event data of the source and target contexts has a specific position in the created feature vectors.
@@ 8.3 Future research
There are four promising directions for future research. First, fine-tuning of the embedding models used in our technique for feature-based TL can be investigated. We suspect that this will improve the prediction performance of the model used in the target context. As the source context can comprise multiple business processes and their event data, certain strategies for fine-tuning can be realized to adapt the embedding models to multiple similar processes in the same context in a controlled way. For example, such a strategy could be to randomly select 1,000 traces from each business process’s event log to ensure a fair fine-tuning across multiple business processes with varying amounts of event data. Second, the transfer of information about additional event log dimensions can be investigated. In this paper, we have considered information on the control flow and the time perspectives in the design of our technique. However, information on additional attributes available in the event data, such as data attributes, might be valuable for the transfer in PPM. For this purpose, approaches from the research stream of multidimensional process representation creation and learning in BPM (Weinzierl et al. 2024) can be helpful to find a representation of multiple event log attributes, facilitating the process knowledge transfer from source to target contexts. The question of which information is necessary for the knowledge transfer is closely connected to the information used and its representation. This question can be investigated with explainable artificial intelligence (XAI) approaches, which are common in PPM (Stierle et al. 2021).
@@
Third, the application of TL in the context of different types of real business processes can be investigated. In the case of those business processes, the complexity of TL is further increased because they differ more structurally and semantically. For example, most of the activities, as well as additional data attributes in the event data of a business process in the source context, can differ from the event data of a business process in the target context.
@@
Fourth, the building of process-specific foundation models that are universally usable for various data-driven BPM tasks along the BPM lifecycle (Rizk et al. 2023; Buss et al. 2024) is another promising avenue for future research. For this, existing foundation models for (multivariate) time series can serve as a starting point for model development (e.g., Ye et al. 2024), while available benchmark event logs and other process-relevant resources (e.g., fundamental BPM books, scientific BPM papers, or additional descriptions of event logs) can be used as a data basis for initial model training. As foundational models can support multiple modalities, types of process data going beyond text-transformed event data (e.g., process model collections or process-attached images, audio tracks, or video sequences) are additionally conceivable for model fine-tuning. In order to fine-tune a foundation model with event data from various organizations, the concept of federated learning can be essential to overcome data privacy restrictions (Verma et al. 2019).
@@ 9 Conclusion
In this paper, we proposed a TL-based technique for PPM, allowing organizations without suitable event data or other relevant resources to create predictions in running business processes for effective decision support. For this purpose, we adopted the computational design science paradigm (Rai et al. 2017) and designed an artifact in the form of a TL-based technique for PPM, consisting of three phases: i) initial model building on source, ii) transfer base model from source to target, and iii) online application of model to target. Based on the instantiation of our technique in both a real-life intra- and an inter-organizational use case, including event logs for ITSM processes from two organizations, we evaluated its prediction performance. Our results suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context.
@@ Appendix A Prediction performance for complete model compared to model sub-variants
Table 12 shows that the complete model of our proposed TL-based technique clearly outperforms its sub-variants in terms of prediction performance when it is trained on $L_{S}$ and tested on $L_{T}$ for the inter-organizational use case. This confirms the importance of using both components in the prediction model for an effective process knowledge transfer from source to target contexts.
@@ Appendix B Prediction performance for different prefix encoding approaches
We applied the index-based encoding, the last-3-state encoding, and the aggregation encoding to our proposed TL-based PPM technique for the inter-organizational use case. We also applied these three encoding approaches to two traditional PPM techniques, using LSTM and XGBoost models. We selected these two traditional PPM techniques because the first uses the same LSTM model as in our proposed technique but not in an TL-setting, and the second uses an XGBoost model, which performed relatively well in the first evaluation of our inter-case organizational use case (see Section 7.2.1).
@@
Table 14: Prediction performance for our TL-based PPM technique that encodes activity and timestamp information about a trace separately compared to a baseline technique that encodes activity and timestamp information about a trace jointly. The baseline technique was tested using the four pre-trained contextualized embedding models of our setup for the inter-organizational use case (average and standard deviation over five runs).